#625 closed flight processing (fixed)

EUFAR17/86, flight day 187/2017, Germany

Reported by: asm Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2017/flight_data/germany/EUFAR17_86-2017_187_Bavarian_Forest

Data arrived from NERC-ARF via hard drive 2017-07-20

Scientific objective: Flight for EUFAR RS4forestEBV Training Course

PI: Andrew Skidmore

Change History (33)

comment:1 Changed 21 months ago by asm

Unpacking

GPS failed for owl flightline 1 (as noted in the logsheet). It does not have a nav file and the time in hdr file is not a gps time (also no darkframes recorded for this file).

Apart from that, the checks raised a few other things to have in mind:

  • The same file, owl 1, missing the dark frames.
  • Owl files 2,3,4 and 28 have one incident with 5,3,6 and 2 dropped frames respectively
  • The last timestamp for owl 18 appears to early
  • Same for fenix 19
  • Incorrect last timestamp for fenix 18
  • 3 empty files for RawLidar: LDR170706_105631_0000000015.scn, /LDR170706_103128_0000000014.scn, LDR170706_101903_0000000015.scn

comment:2 Changed 20 months ago by dac

Basestation

NERC GEF retrieved basestation height from data using Leica GeoOffice: 1.76 m.

comment:3 Changed 20 months ago by dac

Digital Camera

Converted raw images to tif.

comment:4 Changed 19 months ago by asm

Navigation processing

Started. Converted basestation data using: teqc -st 170706064200 -e 170706113500 +nav teqc_output.17n Default_9488_0815_075927.m01 > teqc_output.17o

Done Basestation verification:
Latitude : 48 54 07.72527
Longitude 12 30 46.31861
Ell. Height: 365.196 m
Ant. Height: 1.760 m

comment:5 Changed 19 months ago by asm

Navigation processing

Done. Found a good solution with satellite elevation mask 8 degrees.

comment:6 Changed 19 months ago by asm

Hyperspectral Processing

Started. Some problems with SCTs found along lines. Will process thermal data to see if there was a problem with the Fenix sensor.

comment:7 Changed 19 months ago by asm

Hyperspectral Processing

Logsheet noted that there were some problems with Fenix RScube software. The fps recorded in the hdr files can't produce good geocorrected data. After a few tests, the best fps is 31.95

comment:8 Changed 19 months ago by asm

Hyperspectral Processing

In order to use the new fps, had to set use_nav = false. Found SCT values:

Flightline FENIX
1 8
2 2
3 8
4 4
5 3
6 3
7 2
8 6
9 5
10 3
11 3
12 4
13 3
14 8
15 4
16 4
17 3
18 3
19 2
20 7
21 3
22 3
23 10
24 3
25 6
26 3
27 3
28 3
29 10
30 3

comment:9 Changed 19 months ago by asm

Thermal processing

GPS failed for first flightline. Added GPS times and position to the hdr from Fenix line 1. Also note that there is no navigation file or calibration file for this flightline. Used calibration from flightline 2 to create the stitched file and be able to process the line.

There was no calibration for flightline 8, used the closest in time (which is flightline 7)

Found SCTs.:

Flightline OWL
1 12
2 2
3 2
4 2
5 2
6 2
7 2
8 2
9 2
10 2
11 2
12 3
13 2
14 2
15 2
16 3
17 2
18 2
19 2
20 2
21 2
22 2
23 2
24 2
25 2
26 2
27 2
28 2
29 2
30 2

Please note that when processing flightline 1 need to set the use of navigation to false and for the rest; true

Last edited 19 months ago by asm (previous) (diff)

comment:10 Changed 19 months ago by asm

Hyperspectral Processing

Delivery and readme created. Project is ready for Delivery Check.

comment:11 Changed 19 months ago by asm

Thermal Processing

Delivery and readme created. Project is ready for Delivery Check.

comment:12 Changed 19 months ago by dac

Hyperspectral Delivery Check
Checked delivery a couple of changes needed:

  • Changes needed to logsheet
    • Aircraft base is Straubing Airport
    • GPS Base Station was NERC-ARF
    • Co Pilot was Oliver Smith (I think, have used for rest of logheets)
    • Remove LiDAR as it failed
    • Change from ES17_126-2017_187-RS4forestEBV to ES17_126-2017_187-Bavairan_Forest to match with folder name (or change folder name)
  • Screenshots aren't named correctly - have '.bil' in the name
  • In readme change:
    • "Please note that cloudy weather conditions during flight have caused that the hyperspectral data contains gaps, cloud shadows and a higher than usual number of bands present underflows." to "Weather conditions during flight were cloudy and the data contains a number of pixels which are affected by cloud cover or cloud shadow." The underflows will be due to cloud shadow so I don't think you need to specifically mention
    • "During data acquisition, the framerate was not recorded correctly due to a problem with the sensor leading to a 'streatching' of the mapped data and inaccurate geocorrection. To correct for this the framerate has been manualy etimated, which has enabled a good geogorrection to be optained, validated using Open Street Map vector data. However, small offsets may still remain in specific regions"
  • Check SCT values for 23, 9, 15

comment:13 Changed 19 months ago by dac

Hyperspectral Delivery

  • Updated logsheet to fix problems noted
  • Fixed names for screenshots
  • Corrected my spelling mistakes in suggested updates to readme text

comment:14 Changed 19 months ago by dac

Owl delivery check

Replaced logfile with updated one. Other than that ready to deliver.

Started zipping mapped files.

comment:15 Changed 19 months ago by dac

Owl delivery

Data dispatched via FTP (slot 10).

comment:16 Changed 19 months ago by dac

Fenix delivery check

Aser copied correct navigation files to delivery directory. Ran check_apl_cmd again and outputs look fine.

Marking as ready to deliver. Started zipping mapped files.

comment:17 Changed 19 months ago by asm

Fenix Delivery

Mapped files have been zipped successfully. Have dispatched data via FTP using slot 10.

comment:18 Changed 18 months ago by dac

Digital Camera

Generated delivery. Needs new data quality report then ready for checking.

comment:19 Changed 18 months ago by dac

Digital Camera

Added data quality report. Ready for checking.

comment:20 Changed 18 months ago by asm

Digital camera Delivery check

Started.

comment:21 Changed 18 months ago by asm

Digital Camera Delivery Check

KML file and event file look good.

There is a number of photographs with no usable data due to clouds or saturation that we usually do not deliver. As a resume:
-Saturated or cloud covered 143, 209, 339, 397, 542, 544, 545, 552, 553 (usually removed)
-Medium saturation or partial cloud coverage 142, 337, 396, 554 (can be delivered but please check)
-Low saturation 479 (can be delivered)

Also some photographs are outside the area of coverage (1-5 & 1026 – 1048) but could be useful. Photographs 1036 to 1047 are likely to have been taken during landing as they have the shadow of the plane in the image.

Last edited 18 months ago by asm (previous) (diff)

comment:22 Changed 18 months ago by dac

Digital Camera

Removed images with excessive cloud cover and renumbered files. Kept images outside survey area (apart from ones as plane was landing) and added note to readme. If changes are OK can be delivered.

comment:23 Changed 18 months ago by asm

Digital Camera DC

All checks made again. Project is ready to be delivered.

comment:24 Changed 18 months ago by asm

Digital Camera Delivery

Data dispatched via FTP (slot 10) on 18/10/2017

comment:25 Changed 18 months ago by asm

Data Delivery

Upon request of PI, sent a copy of the Camera, Owl and Fenix data in a 2Tb hard drive on 30/10/1986.

comment:26 Changed 17 months ago by asm

Archiving

Tidied up for archiving.

comment:27 Changed 17 months ago by asm

Hyperspectral Atmospheric Correction

Previously started but not noted in the ticket. In progress.

comment:28 Changed 17 months ago by asm

Hyperspectral Atmospheric Correction

Delivery and readme created. Checked spectra which look sensible for overlapping pixels in different flightlines and created mosaics. Started zipping files for delivery.

Last edited 17 months ago by asm (previous) (diff)

comment:29 Changed 17 months ago by asm

Hyperspectral Atmospheric Corrected Delivery

Files have been zipped correctly. Placed data on FTP slot 10 and also sent a hard drive (22-11-2017).

comment:30 Changed 16 months ago by dac

Archiving

Have created a combined hyperspectral delivery by hardlinking 'EUFAR17_86-187-hyperspectral-20170927' and 'EUFAR17_86-187-hyperspectral-atmospheric-correction-20171115'

comment:31 Changed 16 months ago by dac

Archiving

Tidied project and started uploading files to CEDA.

comment:32 Changed 16 months ago by dac

Archiving

Data uploaded to CEDA.

comment:33 Changed 14 months ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.